# ReUseIt: Synthesizing Reusable AI Agent Workflows for Web Automation

### v1

@@ 7. Discussion
Based on our evaluation findings, here we discuss the implications for design, opportunities with ReUseIt , limitations, and future directions.
@@ 7.1. Implications for Design
Accurate issue detection and explanation lay the foundation for user understanding and trust. Based on our user study results, accurate agent error identification is critical for user understanding and trust of web agents. If the agent misreports its own behavior, user trust can collapse (S2-P2, S2-P5). This feedback aligns with broader research on AI transparency and trust, where users need to monitor and understand the AI’s workings in order to trust it (Kassick, 2024). When an AI system conceals errors or provides only vague error information, users cannot calibrate their trust appropriately and may become frustrated (Kassick, 2024; Berkowitz, 2025).
@@
To boost user trust, systems can incorporate verification mechanisms that confirm an agent’s claims against reality by performing explicit action checks using multiple signals, such as DOM state changes, visual screenshots, and console logs. This type of cross-validation is related to the verifier component in advanced GUI agents, which inspects the environment before and after each action to evaluate whether the result matches the agent’s expectation (Tang et al., 2025b). Along with this, intelligent error discrimination has been explored to filter noise, such as distinguishing transient or benign issues (e.g., a momentary delay in loading) from true failures, suppressing false alarms (Jackson, 2025). We also observed the LLM judge tending to produce false negatives in our experiments (Section 6.2), and these verification techniques can be helpful to ensure a more reliable error detection.
@@
Hierarchical explanations and progressive disclosure of agent messages for better user understanding. Participants in our study emphasized the need for hierarchical explanations to improve a web agent’s interpretability. They wanted the agent’s message delivery to start with a concise top-line summary, followed by more detailed reasoning revealed only on demand (S2-P7). This layered format aligns with the UX principle of progressive disclosure, which suggests showing only essential information first and uncovering complexity progressively (Nielsen, 2006). Such progressive disclosure can help users grasp the outcome without being overwhelmed (Nielsen, 2006; Mueller et al., 2021; Kreiterling, 2025). Furthermore, one participant (S2-P9) specifically endorsed that the agent’s response should clearly identify the problem, explain the cause, and provide guidance on solutions. This format is supported by user-interface guidelines on effective messaging in UX design (Microsoft, 2025b; Minhas, 2024), and can help future research to structure agent messaging in human-agent interaction.
@@
Reveal the process of agent decision-making for debugging and trust. Our participants suggested that visibility into the agent’s decision process can be a helpful strategy to debug failures and build trust. This means exposing why the agent took a certain action and what it considered along the way (Eriksen, 2025). They specifically mentioned that a useful system can reveal the agent’s consideration set over time, e.g., what the agent looked at or weighed before acting (S2-P1). This insight aligns with developments in XAI that use attention mechanisms to make AI’s attention visible (An and Joe, 2022). Translating this to web agents suggests that we could visually indicate which UI elements the agent considered at each step by highlighting them. In this way, users can have the clue to understand how an action was proposed and trace the cause of a failure, such as overlooking a critical button or misprioritizing a less relevant link.
@@
Additionally, another interesting insight was revealing the agent’s inactions (S2-P5). This idea resonates with the concept of contrastive understanding in XAI, which deals with “why”, “why not” questions, and hypothetical “what if” scenarios by comparing the actual outcome with alternatives (Bao and Zeng, 2024). Prior HCI research has demonstrated the value of this approach by showing that explanations for why an action was taken and why others were not can significantly improve system intelligibility (Krarup et al., 2019). Thus, a design that surfaces the UI elements evaluated and explicitly states the detailed rationale can help to build more interpretable web agents.
@@
Correlate text messages with interface actions for verification. Participants in our study requested an auto-transcriber (S2-P4) and simultaneous commentary while the interface is updating (S2-P7) for easy agent behavior understanding. Prior HCI literature echoed that agents need to explain their actions while acting to allow easy inspection (Lieberman and Selker, 2003). This lets users catch mistakes and understand the outcome of each step, which is crucial for acceptance (e.g., justifying decisions, ensuring safety) (Lieberman and Selker, 2003; Rosenfeld and Richardson, 2019). In fact, recent UX guidelines for AI recommend showing live status indicators as side-panel histories of every API call or action, because exposing the AI’s chain of thought reduces user anxiety about black-box behavior by around 34% (Zhu, 2025).
@@
Beyond simply describing actions, it is vital that claims are verifiable against the visuals. Our users desired to see confirmation of what the agent said it did (S2-P5). This matches findings from voice-assistant studies in which users trust results more with visible evidence (Budiu and Laubheimer, 2018). Likewise, if the web agent says “Scrolled to the bottom of the page,” the interface should visibly highlight the bottom or timestamp that action so users can confirm it. Concretely, every narrated step could tag the time in the activity log and outline the relevant UI element. If the agent’s summary does not match the actual UI state, the system can flag a discrepancy so the user notices.
@@ 7.2. Opportunities in ReUseIt
Tailor content of user notification based on user and task. The right amount and style of the message in a user notification depend on both the user and the task. Our participants commented that for simple tasks (e.g., price search, news search), they preferred reliability over interpretability (S2-P3, S2-P7). For complex tasks (e.g., reading papers, aggregating statistics), on the other hand, our participants desired fine-grained workflows to dive deeply into how each action is taken for rigorous user review (S2-P3). These insights suggest valuable research directions for tailoring agent messaging and control according to user needs. We can surface historical success rates via public data (e.g., a leaderboard), so users can quickly gauge how likely their task may succeed. Based on this information, users can have the option to choose an assurance mode to allow the agent to run in the background and return a compact outcome, or a research mode with fine-grained traces like action logs, UI elements considered, and exportable artifacts for detailed review.
@@ 7.3. Limitations and Future Work
Reusability and generalizability tradeoff. Our current workflow synthesis has been affected by the web agent capability, such that extensive condition checks were synthesized from failed executions to improve web agent reusability. In the meantime, it raises a question regarding the reusability vs. generalizability tradeoff. Intuitively, workflows with fewer restrictions can afford better generalizability as the agent can adapt its detailed actions according to varying scenarios. However, our experiments show that agents following fewer restrictions had more random actions, which downgraded their reliability to reasonably generalize. This means that current web agents still require specific and detailed instructions to achieve the expected behavior. As the agent capabilities advance, future work can explore this tradeoff with varying levels of detail in the synthesized workflows.
User intervention with agent execution checkpointing and branching. ReUseIt allows human-in-the-loop intervention at points where the agent fails after retries, which forms a linear way of user intervention where the agent pauses the action, requests user guidance, and resumes the execution. In addition to this, there are cases where users need to intervene at a previous step because an earlier prerequisite is not satisfied, e.g., the agent cannot submit a form because a required entry has not been filled out yet. To support this type of user intervention, future work can explore using agent frameworks like LangGraph (LangGraph, 2025) or debug tools like AGDebugger (Epperson et al., 2025) for automatic checkpointing. These techniques can save the execution state after each action so the agent or user can rewind to that point, update instructions, and fork a new execution. Based on the idea of checkpointing, future research can also study workflow branching to explore alternative execution paths. Prior work has found that Directed Acyclic Graph (DAG)-based workflows with parallel branches can significantly boost the efficiency of LLM agents (Qiao et al., 2024). With human-in-the-loop agent execution, automatically synthesized workflows can benefit from human experience and knowledge to include branches and propagate that to future workflow synthesis for efficient task execution.
@@ 8. Conclusion
End users and professionals often need to execute web tasks repetitively. Although AI-powered web agents have the potential to automate repetitive tasks, they lack the reliability to execute these tasks without human guidance. We present ReUseIt , an automatic approach to synthesize workflows with execution guards that can guide agents during task execution. The execution guards incorporate condition checks and fallback actions that allow agents to identify and recover from errors for higher execution reliability, as well as help users to understand issues for better interpretability. In a benchmark evaluation on fifteen tasks, we found ReUseIt ’s synthesized workflows significantly increased the average success rate from 24.2% to 70.1%. In a user study with nine users, we observed that the synthesized workflows helped users to easily understand agent failures and boosted their trust and willingness to adopt web agents.
Appendix A Formative Study
@@ A.1. Tasks
Table 4 summarizes the tasks used in the formative study.
Table 4. Web automation tasks in the formative study.
@@
A.2. Interview Questions
@@
A.3. Follow-Up Experiment Results
Table 5 shows the follow-up experiment results for each user.
Table 5. Formative study follow-up experiment per-user results (avg $\pm$ std success rate).
@@ A.3. Follow-Up Experiment Results
S1-P1 45.0% $\pm$ 10.0% 85.0% $\pm$ 10.0% 95.0% $\pm$ 10.0% 75.0% $\pm$ 26.5%
S1-P2 73.3% $\pm$ 11.5% 73.3% $\pm$ 11.5% 60.0% $\pm$ 20.0% 68.9% $\pm$ 7.7%
S1-P3 40.0% $\pm$ 0.0% 60.0% $\pm$ 16.3% 75.0% $\pm$ 25.2% 58.3% $\pm$ 17.6%
S1-P4 86.7% $\pm$ 11.5% 86.7% $\pm$ 11.5% 66.7% $\pm$ 11.5% 80.0% $\pm$ 11.5%
S1-P5 50.0% $\pm$ 20.0% 75.0% $\pm$ 19.1% 85.0% $\pm$ 19.1% 70.0% $\pm$ 18.0%
S1-P6 86.7% $\pm$ 11.5% 93.3% $\pm$ 11.5% 86.7% $\pm$ 11.5% 88.9% $\pm$ 3.8%
Appendix B User Study
B.1. Interview Questions

### v2

@@ 7. Discussion
Based on our evaluation findings, here we discuss the implications for design, opportunities with ReUseIt , current limitations, and future directions.
@@ 7.1. Implications for Design
Accurate issue detection and explanation lay the foundation for user understanding and trust. Our user study results reveal that accurate agent error identification is critical for user understanding and trust of web agents. If the agent misreports its own behavior, user trust can collapse (S2-P2, S2-P5). This finding aligns with broader research on AI transparency and trust, where users need to monitor and understand the AI’s workings in order to trust it (Kassick, 2024). When an AI system conceals errors or provides only vague error information, users cannot calibrate their trust appropriately and may become frustrated (Kassick, 2024; Berkowitz, 2025).
@@
To boost user trust, systems can incorporate verification mechanisms that confirm an agent’s claims against reality by performing explicit action checks using multiple signals, such as document object model (DOM) state changes, visual screenshots, and console logs. This type of cross-validation is related to the verifier component in advanced GUI agents, which inspects the environment before and after each action to evaluate whether the result matches the agent’s expectation (Tang et al., 2025a). Along with this, intelligent error discrimination has been explored to filter noise, such as distinguishing transient or benign issues (e.g., a momentary delay in loading) from true failures, suppressing false alarms (Jackson, 2025). We also observed the LLM judge tending to produce false negatives in our experiments (Section 9), and these verification techniques can be helpful to ensure a more reliable error detection.
@@
Hierarchical explanations and progressive disclosure of agent messages for better user understanding. Participants in our study emphasized the need for hierarchical explanations to improve the web agent’s interpretability. They wanted the agent’s message delivery to start with a concise top-line summary, followed by more detailed reasoning revealed only on demand (S2-P7). This layered format aligns with the UX principle of progressive disclosure, which suggests showing only essential information first and uncovering specificity progressively (Nielsen, 2006). Such progressive disclosure can help users grasp the outcome without being overwhelmed (Nielsen, 2006; Mueller et al., 2021; Kreiterling, 2025). Furthermore, one participant (S2-P9) specifically endorsed that the agent’s response should clearly identify the problem, explain the cause, and provide guidance on solutions. This format is supported by user-interface guidelines on effective messaging in UX design (Microsoft, 2025b; Minhas, 2024), and can help future research to structure agent messaging in human-agent interaction.
@@
Reveal the process of agent decision-making for user debugging and trust. Our participants suggested that visibility into the agent’s decision process can be a helpful strategy to debug failures and build trust. This means exposing why the agent took a certain action and what it considered along the way (Eriksen, 2025). They specifically mentioned that a useful system can reveal the agent’s consideration set over time, e.g., what the agent looked at or weighed before acting (S2-P1). This insight aligns with developments in XAI that visualize AI’s attention (An and Joe, 2022). Translating this to web agents suggests that we could visually indicate which UI elements the agent considered at each step by highlighting them. In this way, users can understand how an action was proposed and trace the cause of a failure, such as overlooking a critical button or misprioritizing a less relevant link.
@@
Additionally, another interesting insight was revealing the agent’s inactions (S2-P5). This idea resonates with the concept of contrastive understanding in XAI, which deals with “why”, “why not” questions, and hypothetical “what if” scenarios by comparing the actual outcome with alternatives (Bao and Zeng, 2024). Prior HCI research has demonstrated the value of this approach by showing explanations of why an action was taken and why others were not can significantly improve system understandability (Krarup et al., 2019). Thus, a design that surfaces the UI elements evaluated and explicitly states the detailed rationale can help to build more interpretable web agents.
@@
Correlate text messages with interface actions for verification. Participants in our study requested an auto-transcriber (S2-P4) and simultaneous commentary while the interface is updating (S2-P7) for easy agent behavior understanding. Prior HCI literature echoed that agents need to explain their actions while acting to allow easy inspection (Lieberman and Selker, 2003). This lets users catch mistakes and understand the outcome of each action, which is crucial for user acceptance (e.g., justifying decisions, ensuring safety) (Lieberman and Selker, 2003; Rosenfeld and Richardson, 2019). In fact, recent UX guidelines for AI recommend showing live status indicators as side-panel histories of every API call or action, because exposing the AI’s chain of thought reduces user anxiety about black-box behavior by around 34% (Zhu, 2025).
@@
Beyond simply describing actions, it is vital that claims are verifiable against the visuals. Our users desired to see confirmation of what the agent said it did (S2-P5). This matches findings from voice-assistant studies in which users trust results more with visible evidence (Budiu and Laubheimer, 2018). Likewise, if the web agent says “I scrolled to the bottom of the page,” the interface should visibly highlight the bottom or timestamp that action so users can confirm it. Concretely, every narrated step could tag the time in the activity log and outline the relevant UI element. If the agent’s summary does not match the actual UI state, the system can flag a discrepancy so the user notices.
@@ 7.2. Opportunities in ReUseIt
Tailor user notification based on user and task. The right content and style of user notification should depend on both the user and the task. Our participants commented that for simple tasks (e.g., price search, news search), they preferred reliability over interpretability (S2-P3, S2-P7). For complex tasks (e.g., reading papers, aggregating statistics), on the other hand, our participants desired fine-grained workflows to dive deeply into how each action is taken for rigorous user review (S2-P3). These insights suggest valuable research directions for tailoring agent messaging and control according to user needs. We can surface historical success rates via public data (e.g., a leaderboard), so users can quickly gauge how likely their task may succeed. Based on this information, users can have the option to choose an assurance mode to allow the agent to run in the background and return a compact outcome, or a research mode with fine-grained traces like action logs, UI elements considered, and exportable artifacts for detailed review.
Extend the workflow synthesis approach to support other types of AI agents. Our work introduces a workflow synthesis approach to augment the reliability and reusability of web agents. The idea of synthesizing execution guards to guide agent reflection and recovery applies to broader AI agents, such as GUI agents (e.g., agents in mobile UIs (Wang et al., 2024a) and computer operating systems (Wu et al., 2024; Xie et al., 2024)) and API-based AI agents (Shen et al., 2024). The workflow synthesis can follow a similar trial-and-error process to derive condition checks and fallback actions. To construct condition checks, GUI agents can use screenshots or system states, and API agents can verify the API call status and responses. Besides fully automated approaches, we may design human-in-the-loop approaches where users and agents can examine and correct suspicious or erroneous states together. In these applications, we can still leverage issues surfaced by failed condition checks to notify users to help them monitor and debug agent workflows so they do not need to manually track the progress, which can be especially helpful when lower-level action traces are difficult to interpret (e.g., API agents).
@@ 7.3. Limitations and Future Work
Level of workflow details, and reusability vs. generalizability tradeoff. Our current workflow synthesis has been affected by the web agent capability, such that extensive condition checks were synthesized from failed executions to improve web agent reusability. In the meantime, it raises a question regarding the reusability vs. generalizability tradeoff. Intuitively, workflows with fewer restrictions can afford better generalizability as the agent is expected to adapt its actions in varying scenarios. However, our experiments show that agents following fewer restrictions (e.g., Magentic-UI Plan) had more random actions, which downgraded their reliability to reasonably generalize. This means that current web agents still require specific and detailed instructions to achieve the expected behavior, which can lead to verbose workflows and increase the user’s burden to offer guidance when the agent fails to follow the workflows. We envision that as agents evolve with improved planning and execution capabilities, they can adhere more strictly to planned actions and exhibit less random behavior. In such scenarios, tightened workflows can be synthesized to guide agents. This would reduce the time required to synthesize condition checks and fallback actions from trial-and-error and minimize the need for user oversight and intervention to recover from low-level but critical mistakes. Instead, high-level workflows consisting of major steps could be employed, allowing user involvement to focus on personalizing workflows for their specific needs. With this approach, future work could further explore the tradeoff between reusability and generalizability, balancing the level of details to satisfy user needs for either execution accuracy or task extendability.
User intervention with agent execution checkpointing and branching. ReUseIt allows human-in-the-loop intervention at points where the agent fails after retries, which forms a linear way of user intervention where the agent pauses the action, requests user guidance, and resumes the execution. In addition to this interaction method, there are cases where users may need to intervene at a previous step because an earlier prerequisite is not satisfied, e.g., the agent cannot submit a form because a required entry has not been filled out yet. To support this type of user intervention, future work can explore using agent frameworks like LangGraph (LangGraph, 2025) or debug tools like AGDebugger (Epperson et al., 2025) for automatic checkpointing. These techniques can save the execution state after each action so the agent or user can rewind to that point, update instructions, and fork a new execution. Based on the idea of checkpointing, future research can also study workflow branching to explore alternative execution paths. Prior work has found that Directed Acyclic Graph (DAG)-based workflows with parallel branches can significantly boost the efficiency of LLM agents (Qiao et al., 2024). With human-in-the-loop agent execution, automatically synthesized workflows can benefit from human experience and knowledge to include branches and propagate that to future workflow synthesis for efficient task execution.
@@ 8. Conclusion
End users and professionals often need to execute web tasks repetitively. Although AI-powered web agents have the potential to automate repetitive tasks, they lack the reliability to execute these tasks without human guidance. We present ReUseIt , an automatic approach to synthesize workflows with execution guards that can guide web agents during task execution. The execution guards incorporate condition checks and fallback actions that allow agents to identify and recover from errors for higher execution reliability, as well as help users to understand issues for better agent interpretability. In a benchmark evaluation on fifteen tasks, we found ReUseIt ’s synthesized workflows increased the average success rate from 24.2% to 70.1%. In a user study with nine users, we observed that the synthesized workflows helped users to easily understand agent failures and boosted their trust and willingness to adopt web agents.
Acknowledgements.
We would like to thank Jianfeng Gao, Michel Galley, Swadheen Shukla, and members of the Microsoft Research Deep Learning group for the brainstorming and discussion sessions that helped us shape our project. We want to thank our study participants for their participation and offering us valuable insights on the evaluation. We also want to thank the anonymous reviewers for their constructive feedback that helped us improve our paper.
Appendix A Preliminary Evaluation Tasks and Results
This section presents the tasks used in the preliminary evaluation and formative study, and the preliminary evaluation results (Section 3.1).
@@ A.1. Tasks
Table 5 summarizes the tasks used in the preliminary evaluation and formative study.
Table 5. Preliminary evaluation and formative study web automation tasks.
@@
A.2. Preliminary Evaluation Results
Table 6 summarizes the mean success rates (with 95% confidence intervals) of each task and across all tasks for each condition in the preliminary evaluation.
Table 6. Preliminary evaluation results. Mean success rates (%) $\uparrow$ with 95% confidence intervals of each task and across tasks.
Flight Housing Pet Task-Only 17.5 [9.5, 25.5] 36.7 [22.3, 51.0] 26.7 [-2.0, 55.4] Task + Success-Traces 40.0 [17.5, 62.5] 80.0 [-6.1, 166.1] 73.3 [16.0, 130.7] Task + Magentic-UI Plan 42.5 [22.5, 62.5] 80.0 [30.3, 129.7] 80.0 [30.3, 129.7] Task + Magentic-UI Plan User 45.0 [34.8, 55.2] 82.2 [63.1, 101.3] 84.4 [65.3, 103.6] Price Publication News Task-Only 30.0 [11.6, 48.4] 26.7 [-2.0, 55.4] 30.0 [11.6, 48.4] Task + Success-Traces 45.0 [-15.2, 105.2] 60.0 [-26.1, 146.1] 70.0 [9.1, 130.9] Task + Magentic-UI Plan 62.5 [20.7, 104.3] 63.3 [25.4, 101.3] 80.0 [35.0, 125.0] Task + Magentic-UI Plan User 73.3 [50.4, 96.2] 71.1 [36.6, 105.6] 85.0 [58.5, 111.5]
Across Tasks
27.9 [21.3, 34.5]
61.4 [44.5, 78.3]
68.1 [52.2, 83.9]
73.5 [57.6, 89.4]
Appendix B Formative Study Materials and Results
This section offers the materials (interview questions) used in the formative study and additional results in the follow-up experiments (Section 3.2).
B.1. Interview Questions
@@
B.2. Follow-Up Experiment Results
Table 7 shows the follow-up experiment results for each user in the formative study.
Table 7. Formative study follow-up experiment per-user results. Mean success rates (%) $\pm$ std.
@@ B.2. Follow-Up Experiment Results
S1-P1 45.0 $\pm$ 10.0 85.0 $\pm$ 10.0 95.0 $\pm$ 10.0 75.0 $\pm$ 26.5
S1-P2 73.3 $\pm$ 11.5 73.3 $\pm$ 11.5 60.0 $\pm$ 20.0 68.9 $\pm$ 7.7
S1-P3 40.0 $\pm$ 0.0 60.0 $\pm$ 16.3 75.0 $\pm$ 25.2 58.3 $\pm$ 17.6
S1-P4 86.7 $\pm$ 11.5 86.7 $\pm$ 11.5 66.7 $\pm$ 11.5 80.0 $\pm$ 11.5
S1-P5 50.0 $\pm$ 20.0 75.0 $\pm$ 19.1 85.0 $\pm$ 19.1 70.0 $\pm$ 18.0
S1-P6 86.7 $\pm$ 11.5 93.3 $\pm$ 11.5 86.7 $\pm$ 11.5 88.9 $\pm$ 3.8
Appendix C ReUseIt Prompts
This section offers the prompts used in ReUseIt to generate variation tasks and synthesize condition checks, fallback actions, and workflows (Section 4).
C.1. Variation Task Generation
You are an expert specializing in generating task variations for web agents. Your goal is to analyze a given user task and produce different task variants according to the given requirements.
TASK TO ANALYZE:
{ORIGINAL TASK}
METHOD:
Analyze the input task and identify its variable components (e.g., form values, categories, websites). Based on this analysis, generate three distinct task variations, each corresponding to a different type of variation, while preserving the original task’s objective and intent. The three required variation types are:
  • Attribute Variation: Modify specific input values (e.g., dates, quantities, names, locations, or other form fields) that would be entered on the same webpage.
  • Category Variation: Modify a high-level option that requires switching a tab, toggle, or category within the same website (e.g., changing ‘‘One-way’’ to ‘‘Round-trip’’ on a flight search site).
  • Website Variation: Change the target website to a different one while keeping the underlying task objective unchanged (e.g., switching from Expedia.com to Google Flights for searching flight tickets).
OUTPUT REQUIREMENTS:
Return the output using the exact structure specified below. The original task must be reproduced verbatim, without any modification. Each generated variation must clearly correspond to its designated variation type and remain realistic and actionable for a web agent.
The output must strictly follow the EXAMPLE OUTPUT FORMAT.
EXAMPLE OUTPUT FORMAT:
  • Original Task: Exact original task text.
  • Attribute Variation Task: Task with modified attribute values on the same webpage.
  • Category Variation Task: Task involving a different tab, toggle, or category on the same website.
  • Website Variation Task: Task performed on a different website with the same objective as the original task.
C.2. Condition Check Synthesis
You are an expert specializing in diagnosing failures and formalizing condition checks from these failures for web agents. Your goal is to analyze failed execution messages and convert them into explicit condition checks that should be met to prevent future failures.
INPUT TO ANALYZE:
{FAILED EXECUTION MESSAGES}
METHOD:
Analyze each failed execution message to determine the underlying reason an action did not succeed (e.g., inactive buttons, missing input fields). Identify cues such as ‘‘failed to,’’, ‘‘didn’t,’’, or ‘‘couldn’t’’ to locate failure points. For each failure, abstract the cause into a clear, verifiable condition that should be checked before or after the corresponding action.
Important Constraint:
When deriving condition checks from a failed action, do not include any concrete or literal values from the original action (e.g., specific text strings, numbers, dates, names, or URLs). Conditions must be written using generic, value-agnostic wording that captures the underlying requirement (e.g., element state, page readiness, input availability) rather than the specific instance that caused the failure. The purpose of each condition is to prevent the same class of failure across different inputs or contexts, not to guard against a single, fixed value from the original action.
OUTPUT REQUIREMENTS:
For every identified failure, generate a rule using the exact syntax specified below. Each rule must clearly state:
  • The relevant action.
  • Whether the check should occur before or after the action.
  • The condition that should be satisfied to avoid the failure.
The output must strictly follow the EXAMPLE OUTPUT FORMAT.
EXAMPLE OUTPUT FORMAT:
  • Action: Concise description of the failed action.
  • Condition Check: Before or after performing {Action}, ensure {Condition} is satisfied.
C.3. Fallback Action Synthesis
You are an expert specializing in synthesizing retry strategies for web agents. Your goal is to analyze failed actions alongside successful executions and derive fallback actions that can help web agents recover from failures.
INPUT TO ANALYZE:
Failed Action Description:
{FAILED ACTION}
Successful Execution Messages:
{SUCCESSFUL EXECUTION MESSAGES}
METHOD:
Compare the failed step with the actions described in the successful execution messages of the same action. Identify the specific UI interactions (e.g., clicking, navigating, typing) that enabled success in those executions. Abstract these actions into a clear, specific retry instruction that can be applied to recover from the failure.
Important Constraint:
When deriving fallback actions from a successful action, do not include any concrete or literal values from the original action (e.g., specific text strings, numbers, dates, names, or URLs). Fallback actions must be written using generic, value-agnostic wording that abstracts the underlying recovery strategy (e.g., satisfy element state, ensure page readiness, check input availability), rather than reproducing a specific successful instance. The purpose of each fallback action is to help web agents recover from the same class of failure across varying inputs, pages, or contexts, not limited to addressing a single, fixed failed action.
OUTPUT REQUIREMENTS:
Generate recovery strategies using the exact syntax specified below. The strategies should describe the retry actions in terms of the concrete UI interactions that led to success.
The output must strictly follow the EXAMPLE OUTPUT FORMAT.
EXAMPLE OUTPUT FORMAT:
  • Action: Concise description of the failed action.
  • Fallback Action: Retry {Action} by performing {Fallback Action}.
C.4. Workflow Synthesis
You are an expert specializing in constructing generalizable workflows for web agents. Your goal is to synthesize an executable workflow by integrating condition checks and fallback actions into a given workflow structure.
INPUT TO ANALYZE:
Workflow Structure:
{WORKFLOW STRUCTURE}
Condition Checks:
{CONDITION CHECKS}
Fallback Actions:
{FALLBACK ACTIONS}
METHOD:
For each relevant step in the given workflow structure, insert the corresponding pre- or post-condition checks immediately before or after the action description of that step. In addition, insert fallback actions after the corresponding condition check. Ensure logical ordering and clarity so the workflow can be executed step-by-step by a web agent.
OUTPUT REQUIREMENTS:
Present the final workflow as a numbered list of steps following the given workflow structure. For steps with condition checks and fallback actions, they must clearly label their components using the following tags:
  • Action: Concise description of the action.
  • Condition Check: The condition to verify before or after the action.
  • Fallback Action: The recovery action to take if the condition is not met.
For steps with no condition checks or fallback actions, they must clearly label their components using only the action tag:
  • Action: Concise description of the action.
The output must strictly follow the EXAMPLE OUTPUT FORMAT.
EXAMPLE OUTPUT FORMAT:
For steps without condition check:
  • Action: Perform {Action}
For steps with pre-condition check only:
  • Condition Check: Ensure {Condition}    Fallback Action: Retry {Fallback Action}    Action: Perform {Action}
For steps with post-condition check only:
  • Action: Perform {Action}    Condition Check: Ensure {Condition}    Fallback Action: Retry {Fallback Action}
For steps with both pre- and post-condition checks:
  • Condition Check: Ensure {Condition}    Fallback Action: Retry {Fallback Action}    Action: Perform {Action}    Condition Check: Ensure {Condition}    Fallback Action: Retry {Fallback Action}
Appendix D Benchmark Evaluation Results
Table 8 presents the mean success rates (%) with 95% confidence intervals of each task and across tasks for each condition in the benchmark evaluation (Section 5).
Table 8. Benchmark evaluation results. Mean success rates (%) $\uparrow$ with 95% confidence intervals of each task and across tasks.
Archive Asus Cars Task-Only 13.3 [-15.4, 42.0] 45.0 [14.5, 75.5] 20.0 [20.0, 20.0] Task + Success-Traces 26.7 [-2.0, 55.4] 80.0 [80.0, 80.0] 30.0 [-97.1, 157.1] Task + Magentic-UI Plan 33.3 [4.6, 62.0] 50.0 [-5.1, 105.1] 50.0 [-77.1, 177.1] Task + ReUseIt Workflow 86.7 [58.0, 115.4] 90.0 [71.6, 108.4] 80.0 [80.0, 80.0] Bandcamp BBC Fda Task-Only 40.0 [-214.1, 294.1] 25.0 [-5.5, 55.5] 15.0 [-15.5, 45.5] Task + Success-Traces 50.0 [-331.2, 431.2] 45.0 [-2.7, 92.7] 35.0 [4.5, 65.5] Task + Magentic-UI Plan 80.0 [80.0, 80.0] 60.0 [15.0, 105.0] 40.0 [14.0, 66.0] Task + ReUseIt Workflow 90.0 [-37.1, 217.1] 80.0 [54.0, 106.0] 55.0 [39.1, 70.9] Forbes Gettyimages Groupon Task-Only 15.0 [-32.7, 62.7] 10.0 [-21.8, 41.8] 40.0 [14.0, 66.0] Task + Success-Traces 30.0 [11.6, 48.4] 30.0 [11.6, 48.4] 60.0 [23.3, 96.7] Task + Magentic-UI Plan 35.0 [-12.7, 82.7] 35.0 [4.5, 65.5] 45.0 [5.0, 85.0] Task + ReUseIt Workflow 45.0 [-2.7, 92.7] 55.0 [39.1, 70.9] 80.0 [54.0, 106.0] Restaurantguru Scribd Smithsonianmag Task-Only 20.0 [-25.0, 65.0] 10.0 [-8.4, 28.4] 35.0 [-12.7, 82.7] Task + Success-Traces 30.0 [11.6, 48.4] 35.0 [19.1, 50.9] 55.0 [7.3, 102.7] Task + Magentic-UI Plan 45.0 [5.0, 85.0] 40.0 [14.0, 66.0] 65.0 [49.1, 80.9] Task + ReUseIt Workflow 50.0 [-16.2, 116.2] 55.0 [24.5, 85.5] 90.0 [58.2, 121.8] Sportskeeda Ticketmaster Tvguide Task-Only 20.0 [-25.0, 65.0] 45.0 [14.5, 75.5] 10.0 [-21.8, 41.8] Task + Success-Traces 40.0 [40.0, 40.0] 45.0 [5.0, 85.0] 30.0 [-1.8, 61.8] Task + Magentic-UI Plan 40.0 [-5.0, 85.0] 55.0 [39.1, 70.9] 55.0 [39.1, 70.9] Task + ReUseIt Workflow 55.0 [15.0, 95.0] 65.0 [49.1, 80.9] 75.0 [27.3, 122.7]
Across Tasks
24.2 [16.9, 31.5]
41.4 [33.2, 49.6]
48.6 [41.4, 55.7]
70.1 [61.0, 79.2]
Appendix E User Study Materials and Results
This section provides the materials (interview questions) used in the user study (Section 6.1) and additional results of the follow-up experiments (Section 9).
E.1. Interview Questions